# Optically Switchable Fluorescence Enhancement at Critical Interparticle Distances

### v1

@@ I Introduction
Semiconducting QDs, particularly CdSe/ZnS CSQDs have gathered immense attention due to their size tunable optical properties [1], high quantum yields and sustainable photostability [2]. Their size-dependent emission, a direct result of quantum confinement, allows for precise control over the emitted light’s wavelength, making them invaluable in display technologies [3]. CdSe/ZnS CSQDs exhibit a broad absorption spectrum that extends from the UV to the visible region [4]. The absorption onset is size-dependent, with smaller QDs absorbing at shorter wavelengths (higher energy) and larger QDs absorbing at longer wavelengths (lower energy) [5]. The absorption spectrum typically shows multiple peaks corresponding to electronic transitions between quantized energy levels (i.e. 1S, 1P, 1D) [6]. On the other hand, the photoluminescence (PL) emission of CdSe/ZnS QDs is narrow and symmetric, with a full width at half maximum (FWHM) typically around 20 – 40 nm [7]. The emission wavelength is tunable by varying the core size, ranging from $\sim$ 500 nm (green) to $\sim$ 650 nm (red) for CdSe cores [8].
Superimposed on CdSe core, the ZnS shell plays a crucial role in passivating the CdSe core, significantly enhancing the quantum yield and photostability of these nanocrystals. ZnS capping results in high quantum yields (QYs) of up to 50 – 80 % which is essential for long-duration applications like biological imaging [9] and QD displays [10]. While the development of QD-based sensors is rapidly advancing with applications ranging from environmental monitoring to medical diagnostics [11, 12], display side has majorly been commercialized providing ultra-high-definition resolutions [13].
@@
AgNPs, with their strong localized surface plasmon resonances (LSPRs), offer a compelling platform for enhancing the fluorescence of nearby QDs [17]. The phenomenon of fluorescence enhancement arises from the interaction between the QD’s excitonic transition and the AgNP’s plasmonic field. This interaction can lead to an increase in the QD’s radiative decay rate, resulting in brighter emission and shorter excited-state lifetimes [18]. However, probing the EM interactions to evaluate the degree of heterogeneity in QD-AgNP coupling on a single particle level is is heavily restricted due to fluorescence blinking that majorly hinders the real-time particle tracking [19]. Moreover, experimental results on ensemble QD measurements report various absorption and scattering profiles due to the interference of multiple parameters such as size distribution, interaction with solvent and clustering effects [20].
@@ II Results
of the employed CSQDs as small dots with radius r = 2 nm (1 nm CdSe core + 1 nm ZnS shell), medium dots with radius r = 3 nm (2 nm core CdSe core + 1 nm ZnS shell) and large dots with radius r = 4nm (3 nm CdSe core + 1 nm ZnS shell). Then, to investigate the EM interaction between an incoming EM wave and a single CSQD, we evaluated the absorption and scattering cross section spectra for the three sizes of single CdSe/ZnS CSQDs. We irradiate the CSQDs by a broadband EM wave of 300 - 800 nm plane wave pulse with a 436 nm peak position and 2 fs pulse length that is injected using total-field/scattered-field (TF/SF) method through Huygen’s box source [22, 23]. Fig. 1 (d) shows the obtained absorption cross section (ACS) spectrum. Three-dimensional quantum confinement of charge carriers in CSQDs increases the effective band gap with decreasing CSQD size. Therefore, the optical absorption shifts to the lower wavelengths (higher energies) with decreasing CSQD size. The left-hand side dominating peak (Fig. 1 (d)) in the ultraviolet region demonstrates an enhanced absorption at higher energies (shorter wavelengths) due to the increasing surface coverage of ZnS shell. Since we employed the constant thickness of ZnS shell (1 nm) for increasing core CdSe radius, the resulting UV absorption peak does not exhibit any energy shift. On the other hand, we observe a stark red shift in longer wavelength absorption peak that is associated to the increasing CdSe core radius. The resulting spectral parameters such as peak points, separation bandwitdth and full width at half maximum (FWHM) values are given in the Table 1. Due to the tunable peak separation bandwidth, we emphasize that using the constant-shell-varying-core approach allows controllable band selection (rejection) for the photodiodes using CSQDs. To further investigate the EM interaction between the incoming irradiation and varying size CdSe/ZnS CSQDs, we obtained scattering cross section (SCS) spectrum.
@@
Based on the obtained ACS and SCS profiles, we investigated the emission spectrum of different sized CSQDs. Due to the Stokes shift, the emission peaks redshift compared to the absorption peaks [24]. By adding Stokes shift values of $\Delta\lambda=15.4,17$ and $17.7$ nm for large, mid-size and small CSQDs respectively in conjunction with the experimental data [25], we obtained the emission intensity of the CdSe/ZnS CSQDs as dipole sources in our FDTD simulation (Fig. 2 (a)). Size dependent dipole emission spectra is provided by assigning Lorentzian distributions with peak points at 494.8, 568.3 and 630.4 nm and full width at half maximum (FWHM) linewidths of 27.4, 33.2 and 35.6 nm in compliance with the previous experimental data [26, 27, 5]. It is worth noting that the resulting spectral width is narrower due to the excitation from the lowest energy state and quantum yield variations are omitted to be able to provide a consistent model. Next, we studied a single particle model that the CdSe/ZnS CSQD is placed in a proximity to spherical AgNP to reveal and use the interparticle EM interactions giving rise to fluorescence enhancement. Fig. 2 (b) shows the schematic illustration of the system of materials and their corresponding positioning in free space. We employed a spherical AgNP with a radius of 10 nm around which the EM field is strongly enhanced due to LSPR. Typically, for visible to near-infrared (400 – 800 nm) excitation, this hotspot region extends from the surface of AgNP up to 20 nm. To that end, we investigated the distance dependence of radiative and non-radiative decay rates in the hotspot region (2 – 20 nm). Fig. 2 (c-e) show the distance dependence of radiative and non-radiative decay rates for increasing CSQD radius respectively. In our free space FDTD simulations, we observed that, at short distances, energy is mostly transferred non-radiatively from the CSQD to the AgNP. Since plasmons decay non-radiatively, this behaviour at short interparticle distances is often related to dissipation in which the energy from plasmon excitation is converted into heat via electron-phonon and phonon-phonon interactions. This process includes Förster Resonant Energy Transfer (FRET) mechanism between the dipole source of CdSe/ZnS fluorescent CSQD as a donor and the spherical AgNP as an acceptor. FRET efficiency is strictly related to the spectral overlap between the donor and the acceptor. In our case, the emission peaks and Lorentzian distribution around them spectrally overlaps with AgNP’s peak around 400 - 450 nm. For larger CSQDs that emits around 630 nm, the spectral overlap is less which causes a lower FRET efficiency. As shown in the decay rate plots (Fig. 2 (c-e)), with the increment of CSQD-AgNP interparticle distance, the CSQD’s radiative decay rate starts to overcome the nonradiative decay rate by Purcell factor enhancement boosting the fluorescence emission. As detailed above, FRET quenching dominates at very close separations while the radiative benefits are outstripping at intermediate separations. This is due to the off-resonance configuration in which there is no complete overlap between AgNP and CdSe/ZnS spectra yielding CSQDs with modest fluorescence enhancement.
@@ III Conclusion and Discussion
After Yee’s algorithm to shift electric and magnetic fields in discretized space and time for the solution of Maxwell’s equations [30], FDTD method has been applied to many plasmonic and photonic systems with the advancement of computer technology that allows for the discretization of space in sub-wavelength dimensions [21]. Pure analytical methods fall short in calculating the near fields in complex scenarios, therefore, here we implement an FDTD simulation with Yee cell (mesh) size that is much smaller (1.4 nm) than the incident wavelength (300 - 800nm) that can overcome complex nanostructures with superimposed materials such as CdSe/ZnS CSQDs. By precisely changing the interparticle distance between individual CSQD and AgNP for each CSQD size, we elucidate the EM interactions giving rise to fluorescence enhancement, FRET and non-radiative decay. Our FDTD model provides for the first time critical QD-AgNP interparticle distances by simultaneously measuring the radiative and non-radiative decay rates. We extracted the critical distance points where the suppression of radiative decay is no longer effective, yielding observable fluorescence enhancement in the absence of any gain mechanism. This behavior underscores a central principle in plasmon-enhanced fluorescence: maximal brightness is achieved not simply by maximizing the coupling strength, but by balancing radiative enhancement against nonradiative quenching. Our findings are consistent with theoretical models and previous experimental studies which have demonstrated that the highest fluorescence enhancements are observed in systems where the emitter is optimally detuned or positioned to favor radiative decay over nonradiative losses [31].
@@ IV Methodology
×$\nabla\times\mathbf{H}$ $=-\epsilon\frac{\partial\mathbf{E}}{\partial t}+\mathbf{J},$ (2)
@@
where $k=\omega/c$ is the wavenumber, $\sigma(\omega)$ is the polarizability of the dot, $c$ is the speed of light and $\epsilon_{0}$ is the permittivity of the free space. Radiative and non-radiative decay rates as function of CSQD-AgNP interparticle spacing are calculated to assess the decay dynamics. CdSe/ZnS CSQDs are modelled as dipole sources emitting at the corresponding emission frequencies (absorption wavelength + Stokes shift). Their decay rate in free space is expressed as

### v2

@@ I Introduction
Semiconducting QDs, particularly CdSe/ZnS CSQDs have gathered immense attention due to their size-tunable optical properties [1], high quantum yields and sustainable photostability [2]. Their size-dependent emission, a direct result of quantum confinement, allows for precise control over the emitted light’s wavelength, making them invaluable in display technologies [3]. CdSe/ZnS CSQDs exhibit a broad absorption spectrum that extends from the UV to the visible region [4]. The absorption onset is size-dependent, with smaller QDs absorbing at shorter wavelengths (higher energy) and larger QDs absorbing at longer wavelengths (lower energy) [5]. The absorption spectrum typically shows multiple peaks corresponding to electronic transitions between quantized energy levels (i.e. 1S, 1P, 1D) [6]. On the other hand, the photoluminescence (PL) emission of CdSe/ZnS QDs is narrow and symmetric, with a full width at half maximum (FWHM) typically around 20 – 40 nm [7]. The emission wavelength is tunable by varying the core size, ranging from $\sim$ 500 nm (green) to $\sim$ 650 nm (red) for CdSe cores [8].
Superimposed on CdSe core, the ZnS shell plays a crucial role in passivating the CdSe core, significantly enhancing the quantum yield and photostability of these nanocrystals. ZnS capping results in high quantum yields (QYs) of up to 50 – 80 %, which is essential for long-duration applications like biological imaging [9] and QD displays [10]. While the development of QD-based sensors is rapidly advancing with applications ranging from environmental monitoring to medical diagnostics [11, 12], display side has majorly been commercialized providing ultra-high-definition resolutions [13].
@@
AgNPs, with their strong localized surface plasmon resonances (LSPRs), offer a compelling platform for enhancing the fluorescence of nearby QDs [17]. The phenomenon of fluorescence enhancement arises from the interaction between the QD’s excitonic transition and the AgNP’s plasmonic field. This interaction can lead to an increase in the QD’s radiative decay rate, resulting in brighter emission and shorter excited-state lifetimes [18]. However, probing the EM interactions to evaluate the degree of heterogeneity in QD-AgNP coupling on a single particle level is heavily restricted due to fluorescence blinking that majorly hinders the real-time particle tracking [19]. Moreover, experimental results on ensemble QD measurements report various absorption and scattering profiles due to the interference of multiple parameters such as size distribution, interaction with solvent and clustering effects [20].
@@ II Results
of the employed CSQDs as small dots with radius r = 2 nm (1 nm CdSe core + 1 nm ZnS shell), medium dots with radius r = 3 nm (2 nm core CdSe core + 1 nm ZnS shell) and large dots with radius r = 4nm (3 nm CdSe core + 1 nm ZnS shell). Then, to investigate the EM interaction between an incoming EM wave and a single CSQD, we evaluated the absorption and scattering cross section spectra for the three sizes of single CdSe/ZnS CSQDs. We irradiate the CSQDs by a broadband EM wave of 300 - 800 nm plane wave pulse with a 436 nm peak position and 2 fs pulse length that is injected using total-field/scattered-field (TF/SF) method through Huygen’s box source [22, 23]. Fig. 1 (d) shows the obtained absorption cross section (ACS) spectrum. Three-dimensional quantum confinement of charge carriers in CSQDs increases the effective band gap with decreasing CSQD size. Therefore, the optical absorption shifts to the lower wavelengths (higher energies) with decreasing CSQD size. The left-hand side dominating peak (Fig. 1 (d)) in the ultraviolet region demonstrates an enhanced absorption at higher energies (shorter wavelengths) due to the increasing surface coverage of ZnS shell. Since we employed the constant thickness of ZnS shell (1 nm) for increasing core CdSe radius, the resulting UV absorption peak does not exhibit any energy shift. On the other hand, we observe a stark red shift in longer wavelength absorption peak that is associated to the increasing CdSe core radius. The resulting spectral parameters such as peak points, separation bandwidth and full width at half maximum (FWHM) values are given in Table 1. Due to the tunable peak separation bandwidth, we emphasize that using the constant-shell-varying-core approach allows controllable band selection (rejection) for the photodiodes using CSQDs. To further investigate the EM interaction between the incoming irradiation and varying size CdSe/ZnS CSQDs, we obtained scattering cross section (SCS) spectrum.
@@
Based on the obtained ACS and SCS profiles, we investigated the emission spectrum of different sized CSQDs. Due to the Stokes shift, the emission peaks redshift compared to the absorption peaks [24]. By adding Stokes shift values of $\Delta\lambda=15.4,17$ and $17.7$ nm for large, mid-size and small CSQDs respectively in conjunction with the experimental data [25], we obtained the emission intensity of the CdSe/ZnS CSQDs as dipole sources in our FDTD simulation (Fig. 2 (a)). Size dependent dipole emission spectra is provided by assigning Lorentzian distributions with peak points at 494.8, 568.3 and 630.4 nm and full width at half maximum (FWHM) linewidths of 27.4, 33.2 and 35.6 nm in compliance with the previous experimental data [26, 27, 5]. It is worth noting that the resulting spectral width is narrower due to the excitation from the lowest energy state and quantum yield variations are omitted to be able to provide a consistent model. Next, we studied a single particle model that the CdSe/ZnS CSQD is placed in a proximity to spherical AgNP to reveal and use the interparticle EM interactions giving rise to fluorescence enhancement. Fig. 2 (b) shows the schematic illustration of the system of materials and their corresponding positioning in free space. We employed a spherical AgNP with a radius of 10 nm around which the EM field is strongly enhanced due to LSPR. Typically, for visible to near-infrared (400 – 800 nm) excitation, this hotspot region extends from the surface of AgNP up to 20 nm. To that end, we investigated the distance dependence of radiative and non-radiative decay rates in the hotspot region (2 – 20 nm). Fig. 2 (c-e) show the distance dependence of radiative and non-radiative decay rates for increasing CSQD radius respectively. In our free space FDTD simulations, we observed that, at short distances, energy is mostly transferred non-radiatively from the CSQD to the AgNP. Since plasmons decay non-radiatively, this behaviour at short interparticle distances is often related to dissipation in which the energy from plasmon excitation is converted into heat via electron-phonon and phonon-phonon interactions. This process includes Förster Resonant Energy Transfer (FRET) mechanism between the dipole source of CdSe/ZnS fluorescent CSQD as a donor and the spherical AgNP as an acceptor. FRET efficiency is strictly related to the spectral overlap between the donor and the acceptor. In our case, the emission peaks and Lorentzian distribution around them spectrally overlaps with AgNP’s peak around 400 - 450 nm. For larger CSQDs that emit around 630 nm, the spectral overlap is less which causes a lower FRET efficiency. As shown in the decay rate plots (Fig. 2 (c-e)), with the increment of CSQD-AgNP interparticle distance, the CSQD’s radiative decay rate starts to overcome the nonradiative decay rate by Purcell factor enhancement boosting the fluorescence emission. As detailed above, FRET quenching dominates at very close separations while the radiative benefits are outstripping at intermediate separations. This is due to the off-resonance configuration in which there is no complete overlap between AgNP and CdSe/ZnS spectra yielding CSQDs with modest fluorescence enhancement.
@@ III Conclusion and Discussion
After Yee’s algorithm to shift electric and magnetic fields in discretized space and time for the solution of Maxwell’s equations [30], FDTD method has been applied to many plasmonic and photonic systems with the advancement of computer technology that allows for the discretization of space in sub-wavelength dimensions [21]. Pure analytical methods fall short in calculating the near fields in complex scenarios, therefore, here we implement an FDTD simulation with Yee cell (mesh) size that is much smaller (1.4 nm) than the incident wavelength (300 - 800nm) that can overcome complex nanostructures with superimposed materials such as CdSe/ZnS CSQDs. By precisely changing the interparticle distance between individual CSQD and AgNP for each CSQD size, we elucidate the EM interactions giving rise to fluorescence enhancement, FRET and non-radiative decay. Our FDTD model provides for the first time critical QD-AgNP interpar by simultaneously measuring the radiative and non-radiative decay rates. We extracted the critical distance points where the suppression of radiative decay is no longer effective, yielding observable fluorescence enhancement in the absence of any gain mechanism. This behavior underscores a central principle in plasmon-enhanced fluorescence: maximal brightness is achieved not simply by maximizing the coupling strength, but by balancing radiative enhancement against nonradiative quenching. Our findings are consistent with theoretical models and previous experimental studies which have demonstrated that the highest fluorescence enhancements are observed in systems where the emitter is optimally detuned or positioned to favor radiative decay over nonradiative losses [31].
@@ IV Methodology
×$\nabla\times\mathbf{H}$ $=\epsilon\frac{\partial\mathbf{E}}{\partial t}+\mathbf{J},$ (2)
@@
where $k=\omega/c$ is the wavenumber, $\alpha(\omega)$ is the polarizability of the dot, $c$ is the speed of light and $\epsilon_{0}$ is the permittivity of the free space. Radiative and non-radiative decay rates as function of CSQD-AgNP interparticle spacing are calculated to assess the decay dynamics. CdSe/ZnS CSQDs are modelled as dipole sources emitting at the corresponding emission frequencies (absorption wavelength + Stokes shift). Their decay rate in free space is expressed as